What is x + 2 7/9= 15 1/8?
Answer by unlockmath(1688)

Hello,
For this x + 2 7/9= 15 1/8
Make the fractions improper fraction:
x + 25/9= 121/8
Subtract 25/9 to get:
x = 121/8-25/9
Get common denominators:
x= 1089/72-200/72
x=889/72 or 12 25/72
Make sense?
RJ
www.math-unlock.com